# Interactive4D: Interactive 4D LiDAR Segmentation

### v1

@@ Abstract
Interactive segmentation has an important role in facilitating the annotation process of future LiDAR datasets. Existing approaches sequentially segment individual objects at each LiDAR scan, repeating the process throughout the entire sequence, which is redundant and ineffective. In this work, we propose interactive 4D segmentation, a new paradigm that allows segmenting multiple objects on multiple LiDAR scans simultaneously, and Interactive4D, the first interactive 4D segmentation model that segments multiple objects on superimposed consecutive LiDAR scans in a single iteration by utilizing the sequential nature of LiDAR data. While performing interactive segmentation, our model leverages the entire space-time volume, leading to more efficient segmentation. Operating on the 4D volume, it directly provides consistent instance IDs over time and also simplifies tracking annotations. Moreover, we show that click simulations are crucial for successful model training on LiDAR point clouds. To this end, we design a click simulation strategy that is better suited for the characteristics of LiDAR data. To demonstrate its accuracy and effectiveness, we evaluate Interactive4D on multiple LiDAR datasets, where Interactive4D achieves a new state-of-the-art by a large margin. Upon acceptance, we will publicly release the code and models at https://vision.rwth-aachen.de/Interactive4D.
@@ I INTRODUCTION
This has spurred research into 3D interactive segmentation [23, 53, 46, 15]. Early efforts primarily focused on indoor point clouds, with initial work [23] framing the task as single-object interactive segmentation. In this setting, annotators segment each object individually by providing positive clicks on the object and negative clicks on other areas, essentially treating it as a binary segmentation problem. Recently, [53] reformulated the task as multi-object interactive segmentation where the annotator segments multiple objects simultaneously. Here, the positive clicks for one object inherently serve as negative clicks for other objects, better utilizing user input and increasing efficiency. Both methods only consider interactively segmenting object instances, i.e., things, while neglecting amorphous regions, i.e., stuff.
@@
Despite the successes on indoor point clouds, interactive segmentation of outdoor LiDAR point clouds remains under-explored. Few recent works [46, 15] have tackled this task, yet they follow the single-object paradigm, which is shown to be less efficient than the multi-object paradigm. Also, they treat each LiDAR scan as an independent entity, ignoring the sequential nature of LiDAR scans (Fig. 1, left). Given that LiDAR sensors operate at high frequencies, successive scans capture overlapping regions. Thus, independently annotating each scan is inefficient, leading to unnecessary annotation effort. Furthermore, the independent annotation of each scan complicates the task of maintaining consistent instance IDs across consecutive scans, which is essential for tracking tasks.
Having realized these limitations, in this work, we apply segmenting everything all at once strategy for LiDAR data and propose interactive 4D segmentation, a new paradigm where the annotator segments multiple objects on multiple LiDAR scans simultaneously. To show the effectiveness of this paradigm, we propose Interactive4D, the first interactive 4D segmentation model that performs multi-object segmentation on superimposed consecutive LiDAR scans for both things and stuff objects. This improves efficiency by enabling multi-object interactive segmentation across the entire 4D space-time LiDAR volume. By working directly on 4D data, Interactive4D inherently ensures consistent instance IDs on superimposed consecutive LiDAR scans (Fig. 1, right). This makes it highly adaptable for tracking tasks, simplifying the annotation process for LiDAR tracking datasets, while also paving the way for future research directions.
In the interactive community [31, 23, 53], simulated clicks are commonly used for both training and testing. Many methods typically mimic a user who always clicks at the center of the largest error region [31, 53, 22], while others select clicks at random [23, 28, 46, 15]. Both approaches have notable limitations with LiDAR data. The former incurs high computational costs, causing slower run-times, while both approaches struggle with the sparsity and size variations between small and large objects—common issues in outdoor scenes—often resulting in poor segmentation quality. To overcome these limitations, we propose a new click simulation strategy for both training and evaluation. It generates enhanced and scale-invariant click simulations by accounting for the sparse nature of LiDAR point clouds, identifying the most relevant areas for clicking, and effectively managing scale variations between small and large objects.
@@ III METHOD
Inspired by the success of attention-based models for interactive segmentation [53, 37], we add key technical modifications to achieve the full potential of such models in LiDAR point clouds and introduce Interactive4D, our interactive 4D segmentation model as depicted in Fig. 2. For clarity, we present the entire process using matrix notation.
@@
Localized Loss. (Fig. 2, ) To train the model, we use a combination of the cross-entropy and the dice loss [36].
@@
Here, $E_{i\rightarrow j}$ is the error region consisting of points misclassified as object $j$ on the ground truth object $i$, $p$, and $q$ are the points within and outside the error region respectively. While BD is effective for single object setup and dense data, under multi-object-LiDAR setup, it becomes computationally intensive due to its pairwise distance calculations which are time-consuming and memory demanding. More recent methods [46, 15] mitigate this issue by adopting fully random clicking, significantly reducing the required computation. However, both methods still suffer from two key limitations: (1) Bias Towards Larger Objects: In multi-object interactive segmentation, the error region must first be identified since errors can exist across various objects. BD selection implicitly determines the error size of the region by switching the $\arg\max$ in Eq. (iv) with $\max$ operation. This approach tends to bias clicks toward larger objects (e.g., buildings) while overlooking smaller ones (e.g., bicycles) (Fig. 3, left). The same issue applies to the random click strategy, as the random distribution naturally overlooks smaller objects, making them underrepresented during training. (2) Non-informative Initial Clicks: In dense data, selecting the point farthest from the boundary is effective, as it often captures the “center” of an error region. However, in sparse LiDAR point clouds, this often leads to misplaced clicks near the periphery of the error region due to surrounding empty space (Fig. 3, right). This occurs because points outside the error region determine the boundary. Complete random clicking faces a similar issue, as the selection is not focused on a specific error region but targets the entire set of errors indiscriminately.
@@
This metric, where ||$\cdot$|| ⋅ | indicates the number of points, and $GT_{i}$ is the ground truth object, balances the proportion of the object belonging to the error region, and the current segmentation accuracy of the ground truth object, indicating potential improvement. This approach prioritizes error regions covering a substantial portion of the object, particularly those with low IoU, preventing smaller objects from being overlooked. This results in a more balanced distribution of clicks during training, allowing the model to learn more comprehensive feature representations. (2) Enhanced Click Selection within Error Regions: Similar to prior works that rely on DB, we aim to select the “center” of the ground truth mask for the initial click, as this often captures the key characteristics of the region. To ensure informative selection in sparse domains, we define the center as the point closest to the object’s centroid, typically capturing the most representative area for a solid segmentation start. However, for refinement clicks—where error regions become small and fragmented across the point cloud—the effectiveness of centroid selection decreases and is challenging for real users to follow. To address this, we propose switching to random point selection within the error region for refinement clicks. Unlike  [46, 15], we only perform the random selection after the error region has been identified. The injection of randomness makes the model more robust to user behavior (Sec. IV-2) by learning diverse features.
@@ IV EXPERIMENTS
Datasets. Interactive segmentation involves two key scenarios: (1) Training the model on the annotated portion of the target dataset to assist in labeling the remaining data, referred to as in-distribution. (2) Training the interactive model on a pre-annotated dataset, then applying it to annotate a different target dataset, known as zero-shot. The first highlights its effectiveness in speeding up annotation when partial labels are already available, while the second scenario evaluates the model’s ability to generalize to new, unseen data distributions. To this end, we use two well-established public LiDAR datasets, SemanticKITTI [4] and nuScenes [6]. For all experiments, we train Interactive4D on the training split of SemanticKITTI. The in-distribution setup is evaluated using the validation split of SemanticKITTI, while the zero-shot performance is evaluated on the validation split of nuScenes.
@@
In-distribution Evaluation. We first test our approach by training and evaluating on the same dataset. As shown in Tab. I, Interactive4D achieves significant improvements over the current state-of-the-art, AGILE3D, in all metrics, across all setups. Our Interactive4D, which processes scenes holistically in 4D (Tab. I, last row), surpasses all methods that operate on isolated 3D scans. Notably, with only 1 click per object, it attains an 87.7% $\overline{\text{IoU}}$, outperforming AGILE3D’s 81.5%. Moreover, Interactive4D reaches an 90% $\overline{\text{IoU}}$, a crucial threshold for annotation quality, with just 2.28 clicks.
We also evaluate our 3D multi-object and 4D models (introduced in Tab. I) in the LPS and 4D-LPS tasks on SemanticKITTI and compare it with state-of-the-art non-interactive method  [52], which does not employ any human interaction. As seen in Tab. II, our interactive models significantly outperform this method, even with 1 click per object, and continue to improve with refinement clicks.
@@ IV-1 Click Simulation Strategy
In Tab. V, we demonstrate the effectiveness of our proposed click simulation strategy (Sec. III) compared to different clicking strategies. As shown in Tab. V-➆, our strategy outperforms all others, particularly ➀ and ➁, which are used in previous interactive methods. With just one click, our approach achieves a better $\overline{\text{IoU}}$ (75.3% vs 70.1%) maintaining strong performance at higher click counts (89.8 $\overline{\text{IoU}}@\overline{\text{10}}$). These results highlight that existing strategies, which perform well in dense domains, do not transfer effectively to sparse LiDAR.
We also assessed the robustness of interactive methods when click simulation strategies differ between training and testing. We used AGILE3D clicking strategy only during the testing of Interactive4D. This resulted in an 87.0 $\overline{\text{IoU}}@\overline{\text{10}}$ , outperforming AGILE3D’s 86.6, even though AGILE3D was trained with the same clicking strategy. Moreover, we used our clicking strategy during the testing of AGILE3D, which demonstrated a bigger drop from 86.6 to 81.2 $\overline{\text{IoU}}@\overline{\text{10}}$ (5.4 points). All of these shows that Interactive4D is more robust against different clicking strategies, and our improvements not only originates from the click simulation but also from our architectural design and training.
@@ IV-4 Number of Superimposed Scans in 4D
In Tables I,II, III, all 4D results are presented with 4 superimposed scans to ensure comparability with AGILE3D, which faces memory limitations beyond 4 scans. Fig. IV-3 shows that superimposing more scans increases the segmentation accuracy of Interactive4D even more, especially at lower click counts.

### v2

@@ Abstract
Interactive segmentation has an important role in facilitating the annotation process of future LiDAR datasets. Existing approaches sequentially segment individual objects at each LiDAR scan, repeating the process throughout the entire sequence, which is redundant and ineffective. In this work, we propose interactive 4D segmentation, a new paradigm that allows segmenting multiple objects on multiple LiDAR scans simultaneously, and Interactive4D, the first interactive 4D segmentation model that segments multiple objects on superimposed consecutive LiDAR scans in a single iteration by utilizing the sequential nature of LiDAR data. While performing interactive segmentation, our model leverages the entire space-time volume, leading to more efficient segmentation. Operating on the 4D volume, it directly provides consistent instance IDs over time and also simplifies tracking annotations. Moreover, we show that click simulations are crucial for successful model training on LiDAR point clouds. To this end, we design a click simulation strategy that is better suited for the characteristics of LiDAR data. To demonstrate its accuracy and effectiveness, we evaluate Interactive4D on multiple LiDAR datasets, where Interactive4D achieves a new state-of-the-art by a large margin. We publicly release the code and models at https://vision.rwth-aachen.de/Interactive4D.
@@ I INTRODUCTION
This has spurred research into 3D interactive segmentation [23, 53, 46, 15]. Early efforts primarily focused on indoor point clouds, with initial work [23] framing the task as single-object interactive segmentation. In this setting, annotators segment each object individually by providing positive clicks on the object and negative clicks on other areas, essentially treating it as a binary segmentation problem. Recently, [53] reformulated the task as multi-object interactive segmentation, where annotators segment multiple objects simultaneously. Here, the positive clicks for one object inherently serve as negative clicks for other objects, better utilizing user input and increasing efficiency. Both methods only consider interactively segmenting object instances, i.e., things, while neglecting amorphous regions, i.e., stuff.
@@
Despite the successes on indoor point clouds, interactive segmentation of outdoor LiDAR point clouds remains under-explored. A few recent works [46, 15] have tackled this task, yet they follow the single-object paradigm, which is shown to be less efficient than the multi-object paradigm. Also, they treat each LiDAR scan as an independent entity, ignoring the sequential nature of LiDAR scans (see Fig. 1, left). Given that LiDAR sensors operate at high frequencies, successive scans capture overlapping regions. Thus, independently annotating each scan is inefficient, leading to unnecessary annotation effort. Furthermore, independently annotating each scan complicates the task of maintaining consistent instance IDs across consecutive scans, which is essential for tracking tasks.
Having realized these limitations, in this work, we apply segmenting everything all at once strategy for LiDAR data and propose interactive 4D segmentation, a new paradigm where the annotator segments multiple objects on multiple LiDAR scans simultaneously. To show the effectiveness of this paradigm, we propose Interactive4D, the first interactive 4D segmentation model that performs multi-object segmentation on superimposed consecutive LiDAR scans for both things and stuff objects. This improves efficiency by enabling multi-object interactive segmentation across the entire 4D space-time LiDAR volume. By working directly on 4D data, Interactive4D inherently ensures consistent instance IDs on superimposed consecutive LiDAR scans (see Fig. 1, right). This makes it highly adaptable for tracking tasks, simplifying the annotation process for LiDAR tracking datasets, while also paving the way for future research directions.
In the interactive community [31, 23, 53], simulated clicks are commonly used for both training and testing. Many methods typically mimic a user who always clicks at the center of the largest error region [31, 53, 22], while others select clicks randomly [23, 28, 46, 15]. Both approaches have notable limitations with LiDAR data. The former incurs high computational costs, resulting slower run-times, while both approaches struggle with the sparsity and size variations between small and large objects—common issues in outdoor scenes—that often lead to poor segmentation quality. To address these limitations, we propose a new click simulation strategy for both training and evaluation. It generates enhanced and scale-invariant click simulations by accounting for the sparse nature of LiDAR point clouds, identifying the most relevant areas for clicking, and effectively managing scale variations between small and large objects.
@@ III METHOD
Inspired by the success of attention-based models for interactive segmentation [53, 37], we add key technical modifications to leverage the full potential of such models in LiDAR point clouds and introduce Interactive4D, our interactive 4D segmentation model as depicted in Fig. 2. For clarity, we present the entire process using matrix notation.
@@
Here, $E_{i\rightarrow j}$ is the error region consisting of points misclassified as object $j$ on the ground truth object $i$, $p$, and $q$ are the points within and outside the error region respectively. While BD is effective for single object setup and dense data, under multi-object-LiDAR setup, it becomes computationally intensive due to its pairwise distance calculations which are time-consuming and memory demanding. More recent methods [46, 15] mitigate this issue by adopting fully random clicking, significantly reducing the required computation. However, both methods still suffer from two key limitations: (1) Bias Towards Larger Objects: In multi-object interactive segmentation, the error region must first be identified since errors can exist across various objects. BD selection implicitly determines the error size of the region by switching the $\arg\max$ in Eq. (iv) with $\max$ operation. This approach tends to bias clicks toward larger objects (e.g., buildings) while overlooking smaller ones (e.g., bicycles) (see Fig. 3, left). The same issue applies to the random click strategy, as the random distribution naturally overlooks smaller objects, making them underrepresented during training. (2) Non-informative Initial Clicks: In dense data, selecting the point farthest from the boundary is effective, as it often captures the “center” of an error region. However, in sparse LiDAR point clouds, this often leads to misplaced clicks near the periphery of the error region due to surrounding empty space (see Fig. 3, right). This occurs because points outside the error region determine the boundary. Random clicking exclusively faces a similar issue, as the selection lacks focus on a specific error region and instead targets the entire set of errors indiscriminately.
@@
This metric, where ||$\cdot$|| ⋅ | indicates the number of points, and $GT_{i}$ is the ground truth object, balances the proportion of the object belonging to the error region, and the current segmentation accuracy of the ground truth object, indicating potential improvement. This approach prioritizes error regions covering a substantial portion of the object, particularly those with low IoU, preventing smaller objects from being overlooked. This results in a more balanced distribution of clicks during training, allowing the model to learn more comprehensive feature representations. (2) Enhanced Click Selection within Error Regions: Similar to prior works that rely on BD, we aim to select the “center” of the ground truth mask for the initial click, as this often captures the key characteristics of the region. To ensure informative selection in sparse domains, we define the center as the point closest to the object’s centroid, typically capturing the most representative area for a solid segmentation start. However, for refinement clicks—where error regions become small and fragmented across the point cloud—the effectiveness of centroid selection decreases and is challenging for real users to follow. To address this, we propose switching to random point selection within the error region for refinement clicks. Unlike [46, 15], we only perform the random selection after the error region has been identified. The injection of randomness makes the model more robust to user behavior (see Sec. IV-2) by learning diverse features.
@@ IV EXPERIMENTS
Datasets. Interactive segmentation involves two key scenarios: (1) Training the model on the annotated portion of the target dataset to assist in labeling the remaining data, referred to as in-distribution. (2) Training the interactive model on a pre-annotated dataset and then applying it to annotate a different target dataset, known as zero-shot. The first scenario highlights its effectiveness in speeding up annotation when partial labels are already available, while the second scenario evaluates the model’s ability to generalize to new, unseen data distributions. To this end, we use two well-established public LiDAR datasets, SemanticKITTI [4] and nuScenes [6]. For all experiments, we train Interactive4D on the training split of SemanticKITTI. The in-distribution setup is evaluated using the validation split of SemanticKITTI, while the zero-shot performance is evaluated on the validation split of nuScenes.
@@
In-distribution Evaluation. We first test our approach by training and evaluating on the same dataset. As shown in Tab. I, Interactive4D achieves significant improvements over the current state-of-the-art, AGILE3D, in all metrics across all setups. Our Interactive4D, which processes scenes holistically in 4D (see Tab. I, last row), surpasses all methods that operate on isolated 3D scans. Notably, with only 1 click per object, it attains an 87.7% $\overline{\text{IoU}}$, outperforming AGILE3D’s 81.5%. Moreover, Interactive4D reaches an 90% $\overline{\text{IoU}}$, a crucial threshold for annotation quality, with just 2.28 clicks.
We also evaluate our 3D multi-object and 4D models (introduced in Tab. I) in the LPS and 4D-LPS tasks on SemanticKITTI and compare it with the state-of-the-art non-interactive method [52], which does not employ any human interaction. As seen in Tab. II, our interactive models significantly outperform this method, even with 1 click per object, and continue to improve with refinement clicks.
@@ IV-1 Click Simulation Strategy
In Tab. V, we demonstrate the effectiveness of our proposed click simulation strategy (see Sec. III) compared to different clicking strategies. As shown in Tab. V-➆, our strategy outperforms all others, particularly ➀ and ➁, which are used in previous interactive methods. With just 1 click, our approach achieves a better $\overline{\text{IoU}}$ (75.3% vs. 70.1%) maintaining strong performance at higher click counts (89.8% $\overline{\text{IoU}}@\overline{\text{10}}$). These results highlight that existing strategies, which perform well in dense domains, do not transfer effectively to sparse LiDAR.
We also assessed the robustness of interactive methods when click simulation strategies differ between training and testing. We used AGILE3D clicking strategy only during the testing of Interactive4D. This resulted in an 87.0% $\overline{\text{IoU}}@\overline{\text{10}}$, outperforming AGILE3D’s 86.6%, even though AGILE3D was trained with the same clicking strategy. Moreover, we used our clicking strategy during the testing of AGILE3D, which demonstrated a bigger drop from 86.6% to 81.2% $\overline{\text{IoU}}@\overline{\text{10}}$ (5.4 points). All of these shows that Interactive4D is more robust against different clicking strategies, and our improvements not only originate from the click simulation but also from our architectural design and training.
@@ IV-4 Number of Superimposed Scans in 4D
In Tables I,II, III, all 4D results are presented with 4 superimposed scans to ensure comparability with AGILE3D, which faces memory limitations beyond 4 scans. Fig. IV-2 shows that superimposing more scans increases the segmentation accuracy of Interactive4D even more, especially at lower click counts.